Can I please have help with this question? Thank you in advance.
A man has marked off an area in his yard where he will plant a garden. Find the dimensions of the rectangular area if the length is 3 feet more than 3 times the width and the area is 396 square feet.

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52781)

.

Let w be the width, in feet;  then the length is  (3w+3) feet.


The area equation is


    w*(3w+3) = 396.


To solve it, cancel the common factor 3 in both sides.  You will get then


    w*(w+1) = 132.


From this point, there are 2 ways to complete the solution.


1)  You may reduce the last equation to the standard form quadratic equation and then solve it 

    using either the quadratic formula or factoring.


2)  Or you may guess MENTALLY two factors of the number 132 that differ by one unit

     132 = 11*12.


By doing this way, you obtain the ANSWER momentarily:


    The width is 11 feet; the length is 3*11 + 3 = 36 feet.

Solved.



Answer by MathTherapy(10552)


Can I please have help with this question? Thank you in advance.
A man has marked off an area in his yard where he will plant a garden. Find the dimensions of the rectangular area if the length is 3 feet more than 3 times the width and the area is 396 square feet.